Convert 0.45 mops3 into brooms3 using the following conversion factors: 12 bucket = 1 brush 0.00567 mop = 1 bucket 1500. brush = 145 broom The answer is 1.3*10^-3 please explain step by step on how you got the answer
Answer #1

12 bucket = 1 brush

1 bucket = (1/12) brush= 0.0833 brush

1500 brush = 145 broom

1 brush = 145 broom / 1500 = 0.0967 broom

Since 0.00567 mop = 1 bucket = 0.0833 brush = 0.0833 * 0.0967 broom = 0.00805 broom

So 1 mop = 0.00805 broom / 0.00567 = 1.42 broom

Take cube on both sides

(1 mops)^3 = (1.42 brooms)^3 = 2.863 brooms^3

1 mops^3 = 2.863 brooms^3

0.45 mops^3 = 0.45 * 2.863 brooms^3 = 1.3 brooms^3

0.45 mops^3 = 1.3 brooms^3